Above what Fe2+ concentration will Fe(OH)2 precipitate from a buffer solution that has a pH of 8.57 ? The ?sp of Fe(OH)2 is 4.87×10−17.
Answer #1

Fecot), Fe2+ 204- Ksp for this reaction will be Ksp = [ Fezt] [ord) and given Ksp = 4.87*10-17 : : [ rezt] [om-j* = 4-87x10-1